A balloon held a volume of 16.0L at a temperature of 300K. When the temperature was changed to 400K, what was the new volume of the balloon?
1
Expert's answer
2020-04-23T13:08:00-0400

V1/T1=V2/T2;

V2=V1*T2/T1=16*400/300=21.33 L

Answer: 21.33 L